OpenStudy (anonymous):

how do i find 1/3 of 16

OpenStudy (anonymous):

please help

zepdrix (zepdrix):

Using long division? Or calculator? ;D

OpenStudy (geometry_hater):

divide 16 by 3

OpenStudy (anonymous):

i dont think thats how u do it...

OpenStudy (anonymous):

how do u do it?

OpenStudy (geometry_hater):

Well its easier if you do it by a calculator

zepdrix (zepdrix):

Yah geometry is correct. Multiplying by 1/3 is the same as dividing by 3 c: simple as that!

zepdrix (zepdrix):

If your teacher wants you to do the long division though, then it's a little trickier :D

OpenStudy (anonymous):

thank you but what would it be as a fraction

zepdrix (zepdrix):

change 16 to a fraction \[\frac{ 16 }{ 1 }\] then do the multiplication. \[\frac{ 1 }{ 3 } * \frac{ 16 }{ 1 }\]

OpenStudy (anonymous):

its 16/3

zepdrix (zepdrix):

yayyy \c:/

OpenStudy (anonymous):

so thats the answer

zepdrix (zepdrix):

yes, unless your teacher wants it written as a mixed number (whole number + remainder on top of fraction)

OpenStudy (anonymous):

thats what i want but i know hot to do thaat. thanks